assume that a men's weights are normally distributed with a mean given as 172 lbs and a standard deviation given as 29 lbs based on data from the national health survey If one man is selected at random, find the probability that his weight is more than 167lb.
Answer #1

let X be a men's weight

X~

P( X > 167) = P ( )

= P ( Z > -0.17 )

= 1 - P ( Z < - 0.17)

= 1 - 0.4325

= 0.5675